# Vision-Language Models Struggle to Align Entities across Modalities

### v1

@@ Abstract
Cross-modal entity linking refers to the ability to align entities and their attributes across different modalities. While cross-modal entity linking is a fundamental skill needed for real-world applications such as multimodal code generation, fake news detection, or scene understanding, it has not been thoroughly studied in the literature. In this paper, we introduce a new task and benchmark to address this gap. Our benchmark, MATE, consists of 5.5k evaluation instances featuring visual scenes aligned with their textual representations. To evaluate cross-modal entity linking performance, we design a question-answering task that involves retrieving one attribute of an object in one modality based on a unique attribute of that object in another modality. We evaluate state-of-the-art Vision-Language Models (VLMs) and humans on this task, and find that VLMs struggle significantly compared to humans, particularly as the number of objects in the scene increases. Our analysis also shows that, while chain-of-thought prompting can improve VLM performance, models remain far from achieving human-level proficiency. These findings highlight the need for further research in cross-modal entity linking and show that MATE is a strong benchmark to support that progress.
@@
Iñigo Alonso1, Ander Salaberria2, Gorka Azkune2, Jeremy Barnes2, Oier Lopez de Lacalle2, 1Institute for Language, Cognition and Computation, University of Edinburgh 2HiTZ Center - Ixa, University of the Basque Country UPV/EHU, Correspondence: ialonso@ed.ad.uk
@@ 1 Introduction
We can state, thus, that cross-modal entity linking is a basic ability needed to enable further applications of multimodal artificial intelligence systems. However, to the best of our knowledge, no exhaustive and targeted studies can be found in the literature. To fill that gap, in this paper we analyze the capabilities of current Vision-Language Models (VLM) for cross-modal entity linking. Specifically, we build a new multimodal question-answering benchmark called MATE that contains synthetic images of 3D scenes aligned with textual representations of those scenes (Figure 1). We use synthetic images to control for all the variables of the problem and to easily generate different experiments. The task we use to evaluate the ability of models to perform cross-modal entity linking is shown in Figure 1. Given a pointer attribute which is unique in one of the modalities (e.g., the red color in the first question), we ask for a target attribute of that object which is shown in the other modality (e.g., the name of the referred entity). As the modalities we consider are visual and textual, we have image-to-text (question 1 of Figure 1) and text-to-image tasks (question 2), depending on where the pointer and target attributes exist. MATE is specifically designed for evaluating cross-modal entity linking and alignment, a core skill for multimodal tasks.
We evaluate and analyze open- and closed-weight VLMs on MATE and perform further human evaluations. As a result of our experiments, we find that:
@@
While humans achieve almost perfect performance, current VLMs fail to consistently align attributes across modalities out-of-the-box. Furthermore, VLMs are generally worse for the image-to-text variant; humans show a balanced performance. Finally, the performance of VLMs is heavily influenced by the number of objects in the scene and the target attribute requested, whereas human performance is stable across different configurations.
@@
Our dataset, evaluation results, and code are publicly available. [^1] Our dataset, evaluation results, and code are publicly available at https://github.com/AlonsoApp/MATE
@@ 3 Cross-modal Entity Linking
Cross-modal entity linking refers to the ability to understand that an entity described in two different modalities is actually the same entity. For example, in Figure 1, the red cylinder is represented in two ways: visually, with attributes such as color, shape, and size, and textually, as a set of attribute-value pairs. To evaluate whether VLMs possess this ability, we propose a question-answering task. We create 3D scenes containing multiple objects and provide a textual collection of attribute-value pairs for all objects in the scene (see Figure 1). Importantly, while some attributes are shared across modalities, others are exclusively available in one modality. In Figure 1, the color attribute is not included in the textual modality, while the object’s name appears only in text and other attributes (shape, size, and material) are common to both.
@@
Question 1 in Figure 1 is an image-to-text task because the pointer attribute is only represented in the visual modality and the target attribute in the text. Question 2, instead, is text-to-image, where the pointer attribute “name” is only represented in text, and the target attribute is the color, represented only in the image.
This question-answering task is a suitable proxy for testing the ability of cross-modal entity linking, since to achieve high performance, models must link two different representations of the same entity.
@@ 4 Experiments
4.1 Experimental Setup
We conducted our experiments using a variety of open and closed models (see Appendix C for a full list). For open models, we report results for the top-performing model from each family. The families we considered are LLaVA-1.5 Liu et al. (2024b), LLaVA-1.6 Liu et al. (2024a), Molmo Deitke et al. (2024), Llama 3.2 Dubey et al. (2024), and Qwen2-VL Wang et al. (2024). [^2] These VLMs can be paired with different LLM backbones; we only present results for the best-performing variants: LLaVA 1.5 13B, LLaVA 1.6 34B with Nous-Hermes-2-Yi-34B LLM backbone, Molmo-7B-O-0924, Llama-3.2-11B, Qwen2-VL-7B For closed models, we included GPT-4o, Claude 3.5 and Gemini in our evaluation. [^3] Tested versions: gpt-4o-2024-11-20, claude-3-5-sonnet-20241022 and gemini-1.5-flash All open-weight models where evaluated in a computing cluster using four NVIDIA A100 80GB GPUs. The entire set of experiments was estimated to have required approximately 300 GPU hours. For the closed-weight commercial models, we used their respective official APIs.
We tested all models using zero-, one-, and two-shot prompting. As most evaluated models restrict visual input to a single image, we used one image along with two question-answer examples for two-shot prompting. When the target attribute resided in the visual modality, we included a set of possible answer options to help the models align their responses with the expected terminology (see Appendix F.2 for prompt examples).
@@ 4.2 Human Evaluation
We conducted a human evaluation to quantify the extent to which humans are capable of performing the kind of entity-attribute correlations we expect from VLMs. For this purpose, we selected a subset of 384 examples from our benchmark, ensuring a representative distribution of features. [^4] This subset included at least one combination of every target and pointer attribute across examples with all possible object counts from both image-to-text and text-to-image variants. These examples were distributed among 5 participants. Evaluators were presented with the same image and textual prompt used in the VLM evaluations.
@@ 5 Results
We report the exact match accuracy across all experiments, as they involved either extractive question-answering (for textual target attributes) or multiple-choice question-answering (for visual target attributes). We also report results for the two-shot settings due to their better performance and stability across inferences. We consider this setting a better representation of the capabilities of these models in this task (see Appendix D).
Table 1 presents the overall results of the VLMs on our MATE benchmark, divided into image-to-text and text-to-image variants, as well as their average. While the task is straightforward for humans (approaching 100% accuracy), it remains challenging for current VLMs, although their performance is well above random chance. The best-performing open-weight VLM falls 23.2 absolute points behind human accuracy, highlighting significant room for improvement. Even the highest-performing closed-weight commercial VLM falls short of human performance by 14.6 absolute points. This gap is particularly notable given the expectation of near-perfect accuracy on such a fundamental task. Thus, we can conclude that current VLMs, unlike humans, struggle to consistently link representations across different modalities for the same set of objects, which may limit the capacities of those models for more advanced tasks.
Nevertheless, the results reveal a clear progression in VLM capabilities. There is a noticeable trend between model performance, parameter count, and release date, with more recent and larger models consistently outperforming their counterparts (see complete results in Appendix C).
@@ 6 Analysis
For the text-to-image setting, these steps are followed in reverse. In question 2 of Figure 1, for instance, step 2 would involve combining at least two attributes, such as shape and material, or shape and size.
In our analysis, we break down the cross-modal entity linking task into these three subtasks to analyze VLM performance and identify the most challenging aspects for these models.
@@ 6.1 Visual and textual search
In this section, we evaluate whether VLMs can perform entity-attribute alignment within a single modality (steps 1 and 3 as identified by human annotators). This helps determine whether the poor multimodal results stem from a general failure to extract attributes within the same modality or from the challenges of cross-modal reasoning itself, i.e., linking attribute identification (step 2). We conduct question-answering tasks within the same modality for both the pointer and target attributes, which we call image-to-image (where both attributes exist in the visual modality) and text-to-text (where both attributes exist in the textual modality [^7] Due to implementation issues, the tested VLMs require an input image even for the text-to-text experiments. We ran all the experiments twice, using white images and black images as inputs and found that the performance did not vary. Therefore, we report results using a white image for the text-to-text task.). Following the terminology used by Campbell et al. (2024), we are evaluating disjunctive visual and textual searching capabilities separately.
@@
For each case, models were prompted with the task’s objective and provided two solved questions as a two-shot prompting setting. For image-to-image tasks, we included a set of possible answer options to help models align their responses with the terminology of hidden attributes (see Appendix F.1 for examples of prompts).
@@
Table 3 reports the results of unimodal search experiments (image-to-image and text-to-text). All models show significantly stronger performance compared to the cross-modal entity linking (Table 1). In particular, Qwen2-VL and the closed-weights commercial VLMs achieve near-perfect accuracy, although few VLMs still struggle with entity search. In some cases, we observe performance discrepancies between image-to-image and text-to-text tasks. We hypothesize that, despite strong visual capabilities, the LLM backbones of these VLMs play a critical role in determining overall performance, and thus text-to-text performance tends to be higher.
@@
In Figure 3, we plot the performance of VLMs for different numbers of objects in the scenes, reporting the average accuracies for both image-to-image and text-to-text tasks. The results show that model performance is much more stable than in the cross-modal scenario (Figure 2), indicating that the number of objects is not a significant factor when only a single modality is considered. These findings align with Campbell et al. (2024), whose results on disjunctive visual search tasks show that the number of objects does not affect VLM performance when feature interference is low.
@@
Qwen2-VL performs near perfectly in unimodal tasks; however, it struggles with cross-modal ones (Table 1), suggesting that its difficulties in MATE stem from the second step: linking attribute identification, making it a perfect candidate for further analysis. In Section 6.2, we will examine the performance of this model in greater depth.
@@ 6.2 Linking attribute identification
In Figure 4 (a), we analyze how the number of required linking attributes influences the model’s performance. We measure Qwen2-VL’s accuracy, the top-performing open-weight model, on examples where the target entity can be traced with a minimum of one, a combination of two, or three attributes, in addition to the 3D coordinates. Although the 3D coordinates are unique to each entity, using them to correctly identify an object is more complicated than combining other attributes. This can be observed in the last column of the figure, labeled "3dc", where we plot the accuracy for examples where the target object can only be linked using the 3D coordinates.
@@
To shed more light on which type of attribute is most effective as a linking attribute, we measure the model’s accuracy for examples that can be linked using a single attribute and separate it by attribute types. These results can be seen in Figure 4 (b). These attribute types contribute proportionally when combined in pairs or triplets (results for all possible combinations are provided in Appendix E).
Figure 4: (a) Accuracy of Qwen2-VL for examples where the correct object must be identified using one, a combination of two, or three visual attributes (i.e., attributes that can be perceived visually). These attributes are considered in addition to 3D coordinates, which could always uniquely identify objects across modalities but models struggle to interpret. This is evident from the accuracy of the examples that rely solely on 3D coordinates (denoted as 3dc). (b) Accuracy for examples where the correct object can be linked with just one attribute of a single specific type. For example, in the case of material (mat), when the pointed object is the only object in the scene made of rubber or metal. mat: material, sha: shape, col: color, siz: size, 3dc: 3D coordinates.
@@ 6.3 Chain-of-Thought Evaluation
The main experiments require models to solve the problem in a single step, although humans intuitively approach this task as a series of logical steps. For complex problems, however, VLMs and Large Language Models (LLMs) generally benefit from breaking a task into a sequence of simpler subtasks Wei et al. (2022); Chen et al. (2024). This process, known as Chain-of-Thought (CoT), helps guide the final output toward the correct solution of a complex problem. To verify whether poor model performance was due to an inability to handle the task in one step, we conduct a further evaluation of the same task using CoT prompting techniques. Here we prompt the models first to identify the object in the pointer modality that matches the given pointer attribute. Next, they must list this object’s attributes, identify those that distinguish it from other objects in the pointer modality (linking attributes), locate the corresponding object in the target modality using these linking attributes, and finally return its target attribute. We conduct this experiment in a two-shot setting (see Appendix F.3 for the complete prompt).

### v2

@@ Abstract
Cross-modal entity linking refers to the ability to align entities and their attributes across different modalities. While cross-modal entity linking is a fundamental skill needed for real-world applications such as multimodal code generation, fake news detection, or scene understanding, it has not been thoroughly studied in the literature. In this paper, we introduce a new task and benchmark to address this gap. Our benchmark, MATE, consists of 5.5k evaluation instances featuring visual scenes aligned with their textual representations. To evaluate cross-modal entity linking performance, we design a question-answering task that involves retrieving one attribute of an object in one modality based on a unique attribute of that object in another modality. We evaluate state-of-the-art Vision-Language Models (VLMs) and humans on this task, and find that VLMs struggle significantly compared to humans, particularly as the number of objects in the scene increases. Our analysis also shows that, while chain-of-thought prompting can improve VLM performance, models remain far from achieving human-level proficiency. These findings highlight the need for further research in cross-modal entity linking and show that MATE [^1] Our dataset, evaluation results, and code are publicly available at https://github.com/hitz-zentroa/MATE is a strong benchmark to support that progress.
@@
Iñigo Alonso1, Gorka Azkune2, Ander Salaberria2, Jeremy Barnes2, Oier Lopez de Lacalle2 1School of Informatics, The University of Edinburgh 2HiTZ Center - Ixa, University of the Basque Country UPV/EHU Correspondence: ialonso@ed.ad.uk
@@ 1 Introduction
We can state, thus, that cross-modal entity linking is a basic ability needed to enable further applications of multimodal artificial intelligence systems. However, to the best of our knowledge, no exhaustive and targeted studies can be found in the literature. To fill that gap, in this paper we analyze the capabilities of current Vision-Language Models (VLM) for cross-modal entity linking. Specifically, we build a new multimodal question-answering benchmark called MATE that contains synthetic images of 3D scenes aligned with textual representations of those scenes (Figure 1). We use synthetic images as this methodology has a large advantage over real-world data – namely, we can control for all the variables of the problem without interference from confounding factors such as object-attribute recognition or visual disambiguation, ensuring that we evaluate cross-modal entity linking independently. The task we use to evaluate the ability of models to perform cross-modal entity linking is shown in Figure 1. Given a pointer attribute which is unique in one of the modalities (e.g., the red color in the first question), we ask for a target attribute of that object which is shown in the other modality (e.g., the name of the referred entity). As the modalities we consider are visual and textual, we have image-to-text (question 1 of Figure 1) and text-to-image tasks (question 2), depending on where the pointer and target attributes exist. We evaluate and analyze open- and closed-weight VLMs on MATE and perform further human evaluations. As a result of our experiments, we find that:
@@
While humans achieve almost perfect performance, current VLMs fail to consistently align attributes across modalities out-of-the-box. Furthermore, VLMs are generally worse for the image-to-text variant; humans show balanced performance. Finally, the performance of VLMs is heavily influenced by the number of objects in the scene and the target attribute requested, whereas human performance is stable across different configurations.
@@ 3 Cross-modal Entity Linking
Cross-modal entity linking refers to the ability to understand that an entity described in two different modalities is actually the same entity. For example, in Figure 1, the red cylinder is represented in two ways: visually, with attributes such as color, shape, and size, and textually, as a set of attribute-value pairs.
In order to evaluate whether VLMs possess this ability, we propose a question-answering task. We create 3D scenes containing multiple objects and provide a textual collection of attribute-value pairs for all objects in the scene (see Figure 1). Importantly, while some attributes are shared across modalities, others are exclusively available in one modality. In Figure 1, the color attribute is not included in the textual modality, while the object’s name appears only in text and other attributes (shape, size, and material) are common to both.
@@
Question 1 in Figure 1 is an image-to-text task because the pointer attribute is only represented in the visual modality and the target attribute in the text. Question 2, instead, is text-to-image, where the pointer attribute “name” is only represented in text, and the target attribute is the color, represented only in the image. This question-answering task is a suitable proxy for testing the ability of cross-modal entity linking, since to achieve high performance, models must link two different representations of the same entity.
@@ 4 Experiments
We conducted our experiments using a variety of open and closed models (see Appendix D for a full list). For open models, we report results for the top-performing model from each family. The families we considered are LLaVA-1.5 Liu et al. (2024b), LLaVA-1.6 Liu et al. (2024a), Molmo Deitke et al. (2024), Llama 3.2 Dubey et al. (2024), and Qwen2.5-VL Bai et al. (2025). [^2] These VLMs can be paired with different LLM backbones; we only present results for the best-performing variants: LLaVA 1.5 13B, LLaVA 1.6 34B with Nous-Hermes-2-Yi-34B LLM backbone, Molmo-7B-O-0924, Llama-3.2-11B, Qwen2-VL-7B, Qwen2.5-VL-7B For closed models, we included GPT-4o, Claude 3.5 and Gemini in our evaluation. [^3] Tested versions: gpt-4o-2024-11-20, claude-3-5-sonnet-20241022 and gemini-1.5-flash All open-weight models where evaluated in a computing cluster using four NVIDIA A100 80GB GPUs. The entire set of experiments was estimated to have required approximately 300 GPU hours. For the closed-weight commercial models, we used their respective official APIs.
We tested all models using zero-, one-, and two-shot prompting. As most evaluated models restrict visual input to a single image, we used one image along with two question-answer examples for two-shot prompting. When the target attribute resided in the visual modality, we included a set of possible options to help the models align their responses with the expected terminology (see Appendix H.2).
@@ 4.1 Human Evaluation
We furthermore conducted a human evaluation to quantify the extent to which humans are capable of performing the kind of entity-attribute correlations we expect from VLMs, which serves as an upper-bound for the task.
For this purpose, we selected a subset of 384 examples from our benchmark, ensuring a representative distribution of features. [^4] This subset included at least one combination of every target and pointer attribute across examples with all possible object counts from both image-to-text and text-to-image variants. The evaluation examples were distributed among 5 participants (see Appendix C for further details). All of the evaluators were presented with the same image and textual prompts as those used in the VLM evaluations.
@@ 5 Results
We report the exact match accuracy across all experiments, as they involved either extractive question-answering (for textual target attributes) or multiple-choice question-answering (for visual target attributes). We also report results for the two-shot settings due to their better performance and stability across inferences. We consider this setting a better representation of the capabilities of these models in this task (see Appendix F).
Table 1 presents the overall results of the VLMs on our MATE benchmark, divided into image-to-text and text-to-image variants, as well as their average. While the task is straightforward for humans (approaching 100% accuracy), it remains challenging for current VLMs, although their performance is well above random chance. The best-performing open-weight VLM falls 17.8 absolute points behind human accuracy, highlighting significant room for improvement. Even the highest-performing closed-weight commercial VLM falls short of human performance by 14.6 absolute points. This gap is particularly notable given the expectation of near-perfect accuracy on such a fundamental task. Therefore, we can conclude that current VLMs, unlike humans, struggle to consistently link representations across different modalities for the same set of objects, which may limit the capacities of those models for more advanced tasks where this ability is required.
Nevertheless, the results reveal a clear progression in VLM capabilities. There is a noticeable trend between model performance, parameter count, and release date, with more recent and larger models consistently outperforming their counterparts (see complete results in Appendix D).
@@ 6 Analysis
For the text-to-image setting, the same steps are followed, but in reverse order. In question 2 of Figure 1, for instance, step 2 would involve combining at least two attributes, such as shape and material, or shape and size. In our analysis, we break down the cross-modal entity linking task into these three subtasks to analyze VLM performance and identify the most challenging aspects for these models.
@@ 6.1 Visual and textual search
In this section, we evaluate whether VLMs can perform entity-attribute alignment within a single modality (steps 1 and 3 as identified by human annotators). This helps determine whether the poor multimodal results stem from a general failure to extract attributes within the same modality or from the challenges of cross-modal reasoning itself, i.e., linking attribute identification (step 2).
We conduct question-answering tasks within the same modality for both the pointer and target attributes, which we call image-to-image (where both attributes exist in the visual modality) and text-to-text (where both attributes exist in the textual modality). [^7] Due to implementation issues, the tested VLMs require an input image even for the text-to-text experiments. We ran all the experiments twice, using white images and black images as inputs and found that the performance did not vary. Therefore, we report results using a white image for the text-to-text task. Following the terminology used by Campbell et al. (2024), we are evaluating disjunctive visual and textual searching capabilities separately.
@@
For each case, models were prompted with the task’s objective and provided two solved questions as a two-shot prompting setting. For image-to-image tasks, we included a set of possible answer options in order to help the models align their responses with the terminology of hidden attributes (see Appendix H.1 for examples of these unimodal prompts).
@@
Table 3 reports the results of the unimodal search experiments (image-to-image and text-to-text). All models show significantly stronger performance compared to the cross-modal entity linking (Table 1). In particular, Qwen2-VL, Qwen2.5-VL, and the closed-weights commercial VLMs achieve near-perfect accuracy, although a number of VLMs still struggle with entity search. In a number of cases, we observe certain performance discrepancies between image-to-image and text-to-text tasks. We hypothesize that, despite having strong visual capabilities, the LLM backbones of these VLMs play a critical role in determining overall performance, and thus text-to-text performance tends to be higher.
@@
In Figure 5 (see Appendix E), we plot the performance of VLMs for different numbers of objects in the scenes, reporting the average accuracies for both image-to-image and text-to-text tasks. The results show that model performance is much more stable than in the cross-modal scenario (Figure 2), indicating that the number of objects is not a significant factor when only a single modality is considered. These findings align with Campbell et al. (2024), whose results on disjunctive visual search tasks show that the number of objects does not affect VLM performance when feature interference is low.
@@
Qwen2.5-VL performs near perfectly in unimodal tasks; however, it struggles with cross-modal ones (Table 1), suggesting that its difficulties in MATE stem from the second step: linking attribute identification, making it a perfect candidate for further analysis. In Section 6.2, we will examine the performance of this model in greater depth.
@@ 6.2 Linking attribute identification
In Figure 4 (a), we analyze how the number of required linking attributes influences the model’s performance. We measure Qwen2.5-VL’s accuracy, the top-performing open-weight model, on examples where the target entity can be traced with a minimum of one, a combination of two, or three attributes, in addition to the 3D coordinates. Although the 3D coordinates are unique to each entity, using them to correctly identify an object is more complicated than combining other attributes. This can be observed in the last column of the figure, labeled "3dc", where we plot the accuracy for examples where the target object can only be linked using the 3D coordinates.
@@
To shed more light on which type of attribute is most effective as a linking attribute, we measure the model’s accuracy for examples that can be linked using a single attribute and separate it by attribute types. These results can be seen in Figure 4 (b). These attribute types contribute proportionally when combined in pairs or triplets (results for all possible combinations are provided in Appendix G).
These results highlight a clear underperformance of the model in cases where 3D coordinates are the only linking attribute that uniquely identifies the target object. In these cases, the model continues to rely on attribute overlap to make predictions, with a clear preference towards objects with more shared attributes with the target object rather than uniquely relying on the interpretation of 3D coordinates (see Appendix E). However, we still see that the model has a partial understanding of 3D coordinate interpretation. In cases where the predicted object matched all attributes except for the 3D coordinates, the model was able to predict the correct object in 63% of these.
There are still 14.8% of cases where the predicted object also contained a mismatch in one or more non-coordinate attributes, indicating that the failure was not uniquely due to misinterpreting spatial information but also involved other attribute-level errors. Even in these cases, the model selected an object with a lower Euclidean distance to the gold object in 70% of instances, compared to the average distance of objects with the same number of overlapping attributes.
Figure 4: (a) Accuracy of Qwen2.5-VL for examples where the correct object must be identified using one, a combination of two, or three visual attributes (i.e., attributes that can be perceived visually). These attributes are considered in addition to 3D coordinates, which could always uniquely identify objects across modalities but models struggle to interpret. This is evident from the accuracy of the examples that rely solely on 3D coordinates (denoted as 3dc). (b) Accuracy for examples where the correct object can be linked with just one attribute of a single specific type. For example, in the case of material (mat), when the pointed object is the only object in the scene made of rubber or metal. mat: material, sha: shape, col: color, siz: size, 3dc: 3D coordinates.
@@ 6.3 Chain-of-Thought Evaluation
The main experiments require models to solve the problem in a single step, although humans intuitively approach this task as a series of logical steps. For complex problems, however, VLMs and Large Language Models (LLMs) generally benefit from breaking a task into a sequence of simpler subtasks Wei et al. (2022); Chen et al. (2024). This process, known as Chain-of-Thought (CoT), helps guide the final output toward the correct solution of a complex problem.
In order to verify whether poor model performance was due to an inability to handle the task in one step, we conduct a further evaluation of the same task using CoT prompting techniques. Here we prompt the models first to identify the object in the pointer modality that matches the given pointer attribute. Next, they must list this object’s attributes, identify those that distinguish it from other objects in the pointer modality (linking attributes), locate the corresponding object in the target modality using these linking attributes, and finally return its target attribute. We conduct this experiment in a two-shot setting (see Appendix H.3 for the complete prompt).
@@
6.4 Self-Reflection Evaluation
To evaluate whether self-reflective techniques would tackle this problem better, we evaluated the open-weight self-reflecting visual language model VL-Rethinker-7B Wang et al. (2025) using the same 2 shot prompt as models in Table 1. VL-Rethinker-7B achieves 75.7% and 83.4% accuracy in the image-to-data and data-to-image settings, respectively. In comparison, Qwen2.5-VL—its base, non-self-reflective counterpart, achieves 75.7 and 84.5 points in the same settings. This suggests that self-reflecting does not provide any added value for cross-modal entity linking.